(2√7-4√3)(4√+8√3) =??
Answer by stanbon(57419)

(2√7-4√3)(4√7+8√3) =??
------------------
Use FOIL:
F: 8*7 = 56
O: 16sqrt(21)
I: -16sqrt(21)
L: -32*3 = -96
------------------
Combine to get: -40
=============================
Cheers,
Stan H.